OpenStudy (sissyedgar):

Solve 2x + 5 = 27 x =

pooja195 (pooja195):

Subtract 5 from both sides

pooja195 (pooja195):

27-5=?

OpenStudy (sissyedgar):

22... but the total is 11, just figured it out! :P

OpenStudy (anonymous):

x is 10, because 2*10=20 and 20+5=25